In a bag there are 2red,3yellow,4green,and 6blue marbles.What is the probabilty of p(yellow or green)?
Answer by stanbon(75887)

In a bag there are 2red,3yellow,4green,and 6blue marbles.What is the probabilty of p(yellow or green)?
------------------------------
P(yellow or green) = (3+4)/15 = 7/15
=========================================
Cheers,
Stan H.